# Topologically Protected Surface Altermagnetism on Antiferromagnets

### v1

@@ Abstract
Here, we take a different path towards realizing AM: on the surface of conventional AFMs. From a symmetry perspective the surface of a crystal is interesting because bulk symmetries are typically broken at the surface by the loss of translational invariance and any point group symmetries incompatible with the surface normal. This symmetry reduction plays a central role in the emergence of surface states with properties forbidden in the bulk, such as in magnetic and topological systems [10, manchon_new_2015]. Furthermore, in many materials, surface states can benefit from topological protection, such that they are robust against all symmetry-preserving surface perturbations. This raises the enticing possibility of topologically protecting AM surface states in a conventional AFM—which has yet to be demonstrated and is the focus of this investigation.
@@
As a concrete example for surface magnetism, consider the surface of a conventional AFM. Cutting the crystal along a plane such that only a single spin species exists at the surface, one generates a net surface magnetization—a surface ferromagnet with a spin-split surface state [3, 6, 39]. Similarly, an AFM with non-vanishing surface octupolar moment exhibits an altermagnetization [20, 7]. Although the realization of AM at the surface of an AFM was proposed early on, research has focused so far on identifying a spin-split band structure [37, 53, 38]. However, spin-splitting is not a useful signature for surface AM. The characteristic altermagnetic spin density can emerge as a local feature at the surface. Global bulk properties, e.g., the band structure, cannot detect this effect, because the contributions of opposite surfaces often cancels out.
@@
Figure 1: Altermagnetic surface states of spin-degenerate AFMs. In all cases the bulk of the crystal is spin degenerate due to a combined time reversal inversion symmetry. (a) In a trivial AFM the non-altermagnetic bulk symmetry group can break to an altermagnetic surface symmetry group, leading to a boundary Fermi surface which is spin split. (b) In AFM Dirac nodal line semimetals spin-degeneracy is lost at the surface resulting in topologically protected altermagnetic drumhead states. (c) In AFM Dirac semimetals altermagnetic Fermi arcs connect the bulk topological nodes. (d) To unearth the surface behavior described in (a-c) crystals are studied in a slab geometry, i.e., periodic boundary conditions along $x$ and $y$ direction and open boundary conditions in $z$ direction. Here, the crystal structure of Eq. (1) is shown, where spin-up (spin-down) sites are red (blue). The bulk 4-site unit cell is highlighted in blue, labelled by sublattice, and the layer degree of freedom is in the $z$ direction.
@@
In this work, we demonstrate that spin-degenerate AFMs can exhibit altermagnetic spin-split characteristics at the surface. Such a system may be understood as a bulk AFM with local altermagnetic surface states. Note this is substantially different from the surface states of bulk AMs [43, 15, 14, 48, 33, 35, 11, 34] or the influence of an altermagnetic layer on non-magnetic surface states [21]. Of special interest are the cases where these surface states are topologically protected by a bulk symmetry.
@@
We report on both trivial and topologically-protected altermagnetic surface states, providing a minimal model that realizes both of them. First, we explain the general symmetry prerequisites for such surface states to emerge and demonstrate surface AM in a simple, topologically trivial system with spin-degenerate bulk bands and a spin-split surface, see Fig. 1 (a). Next, we consider using topology as a resource to protect the altermagnetic surface. Finding such a protected state is challenging because topology in collinear magnets is typically induced by SOC [Nagaosa10, Bernevig22], however AM is not well-defined in the presence of SOC [47]. Despite the symmetry conditions for band topology and surface AM being close to incompatible, we find that antiferromagnetic nodal line and Dirac semimetals can exhibit topologically-protected altermagnetic surface states. We study two examples; a weak topological phase where the surface is characterized by altermagnetic drumhead states, depicted in Fig. 1 (b), and a Dirac semimetal hosting altermagnetic Fermi arcs, depicted in Fig. 1 (c). Finally, we show that the room-temperature antiferromagnetic Dirac semimetals CuMnAs and CuMnP [36], known for their applications in spintronics [49, 45, 44], are existing material realization of our theory.
@@
Symmetry requirements—Surface AM arises in bulk collinear AFMs when the non-altermagnetic bulk symmetry group breaks down at the surface to an altermagnetic symmetry group. A bulk AFM has a compensated spin-degenerate band structure, however the magnetic ordering induces a compensated even-wave spin-splitting in the spectral density at the surface. We start by providing the exact symmetry requirements for surface AM effect to occur.
@@
Without loss of generality we assume a (001) surface. Hence, there are two symmetries in a collinear magnet which protect spin-degeneracy in the bulk, but are absent at the surface: (i) an inversion $\mathcal{I}$ times time reversal symmetry $\mathcal{T}\parallel\mathcal{IT}]$ and (ii) a $C_{2}$ rotation in spin space, effectively acting as a spin inversion, times a fractional, (intra-unitcell) translation that involves some translation $\bm{t}_{z}$ in the $z$-direction $C_{2}\parallel\bm{t}_{z}\bm{t}_{\text{any}}]$—here $\bm{t}_{\text{any}}$ indicates that the translation can have any component in the $x$ or $y$ direction. At least one of these two symmetries has to be present, but there must not be any other symmetry that protects spin degeneracy and is robust at the surface, e.g., $C_{2}\parallel\bm{t}_{\text{x or y}}]$. Additionally, there must not be any $\mathcal{T}\parallel C_{2z}\mathcal{T}]$ axis, which would forbid spin splitting at the $z=0$ plane. Finally, one combined spatial and time-reversal symmetry is required to enforce a vanishing magnetic moment at the surface, e.g. $\mathcal{T}\parallel\mathcal{T}C_{4z}]$ and $\mathcal{T}\parallel\mathcal{T}M]$ where $M$ is a mirror. The surface state is then altermagnetic, up to known exceptions [30].
@@
Figure 2: Trivial altermagnetic surface state in a spin degenerate AFM, defined in Eq. 1. (a) The 3D bulk band structure is spin-degenerate. (b-e) Colorplots of the spin-resolved spectral function with the following color scheme: $\mathcal{A}_{{\uparrow}}-\mathcal{A}_{{\downarrow}}$ sets the coloring from spin-down polarized (blue) over spin-degenerate (black) to spin-up polarized (red); $\mathcal{A}_{{\uparrow}}+\mathcal{A}_{{\downarrow}}$ sets the opacity. (b,c) The bulk spin spectral density in the slab geometry of height $L_{z}=30$ along a high symmetry path (c) and at the Fermi energy ($E=0$, dashed line) (d). The bulk spectral density is found projecting onto the layer at $z=15$. (d,e) The spin spectral density in the slab geometry at the surface along a high symmetry path (c) and at the Fermi energy (d). (f) The layer-resolved spin conductivity, expressed as the spin splitter angle $\alpha=2\arctan((\sigma^{\uparrow}_{xx}-\sigma^{\downarrow}_{xx})/(\sigma^{\uparrow}_{xx}+\sigma^{\downarrow}_{xx}))$ shows that the the spin splitting is only sizable at the surface.
@@
where $m$ is the local magnetisation that leads to the formation of Néel ordering, $s={\uparrow},{\downarrow}$ is the spin, $t=1$ parametrises the nearest neighbour hoppings, $t^{\prime}$ controls the strength of the next-nearest neighbour hoppings within each layer, and $\delta$ determines the degree of anisotropy in the NNN hoppings. These hoppings are shown explicitly in Fig. 1 (d). Here, we consider only this non-interacting Hamiltonian, however Eq. (1) may be straightforwardly derived as mean-field description of a Hubbard Hamiltonian with sub-leading NNN couplings, see SM.
@@
Individually, each $xy$-layer of this model would be $d$-wave altermagnetic, because its magnetic sublattices are related by a $\pi/2$-rotation. This is simply the result of the checkerboard square lattice structure of each layer, referred to as the DLKK model [9, 27, 28] or $tt^{\prime}-$$\delta$-Hubbard model [9, 17]. However, neighboring layers have inverted spin patterns, ensuring that the spin splitting on odd layers perfectly compensates that on even layers. Thus, the overall magnetic state forms a bulk AFM. Formally, the spin up and spin down sublattices are related by a $z$ intra-unitcell translation $\bm{t}_{z}$, and inversion $\mathcal{I}$, which maps adjacent layers onto one another. In spin-space group notation the symmetries $\mathcal{T}\parallel\mathcal{I}]$ and $C_{2}\parallel\bm{t}_{z}]$ protect the spin degeneracy of the band structure, see Fig. 2 (a).
@@
At a (001) surface the inversion symmetry and $z$ translation symmetry are broken. Hence, surface states in this antiferromagnet are not locally protected against spin splitting. To see this, we consider Eq. 1 in a slab geometry, i.e., a lattice of $L_{z}$ $x$-$y$-layers with open boundary conditions in the $z$-direction and periodic boundary conditions in the $x$- and $y$-directions. Retaining $x$-$y$ translation symmetry allows us to transform to momentum space $\bm{k}=(k_{x},k_{y})$, such that the $L_{z}$ layers appear as $4L_{z}$ different bands in the slab-geometry band structure, see Fig. 2 (b). This model is physically motivated from a Hubbard model—from unrestricted mean-field simulations we found that the magnetization in the 3D slab system is roughly constant throughout the system, see SM.
@@
Surprisingly, even the slab band structure is spin degenerate. This is due to a global inversion plus time-reversal symmetry $\mathcal{T}\parallel\mathcal{I}]$, which forces $\bm{k}$-momentum states with opposite spin at opposite sides of the system to have the same energy iii In fact the global inversion symmetry only exists if there is an even number of layers. For an odd number of layers it is weakly broken but this is an $\frac{1}{L_{z}}$ effect and means that the dispersion of the surface states on opposite sides of the system is slightly different.. However, this has deeper implications: even though spin up and down $\bm{k}$-momentum states are symmetry-constrained to have the same energy they are highly non-local. Due to inversion, their support is at opposite sides of the system. Hence, we can observe a spin-split surface state on one side of the system and its spin-inverted counterpart on the other side of the system.
@@
We visualize the spin-, momentum-, and space-resolved spectral density by the spin-resolved spectral function, which may be written in terms of the Greens function, $G(\omega,\bm{k})$ as
@@
Finally, we also computed the layer-resolved (again we consider pairs of layers) spin conductance, see SM for details. Here, we use the spin-splitter angle $\alpha$ to quantify the spin-splitting, i.e., the angle between the spin-up and down current which is determined via the ratio of spin over electric conductivity, [16]. As expected from the spin-split surface state, the spin conductance is only non-zero at the surface and odd under global inversion, see Fig. 2 (f).
@@
Figure 3: Topological protected altermagnetic surface states in an AFM, described by $H_{\text{topo}}$ following Eqs. 3 and 4. Subplots (b-k) are calculated on a slab geometry with height $L_{z}=31$. (a) The 3D bulk band structure of the Dirac nodal line semimetal (black, $t^{\prime}=0$) and the Dirac semimetal (yellow, $t^{\prime}=0.2$) is spin degenerate. Dashed lines indicate overlap of both band structures. The touching points at E=0.1E=-0.1 (shifted with a small chemical potential $\mu$) are the nodal line and the Dirac points. The spin-resolved spectral function in a slab geometry of the Dirac nodal line semimetal is shown in (b-e) and the Dirac semimetal in (f-i). The color scheme is identical to Fig. 2. Panel (b,f) and (d,h) show the bulk and surface spin spectral density of each model along a high symmetry path respectively. The bulk ($z=15$) spin spectral density at the Fermi energy ($E=0$, dashed line) shows the Dirac nodal line (c) and the Dirac cones (g). The surface ($z=0$) spin spectral density at the Fermi energy shows the corresponding surface states; closed contours where the drumhead states cross the Fermi energy (e) and open arcs where the Fermi arcs cross the Fermi energy (i). (j,k) The layer-resolved spin-splitter angles $\alpha_{x}$ ($\alpha_{y}$), as solid (dashed) line, which can be observed when applying an electric field in $x$-direction ($y$-direction). Panels (j) and (k) correspond to the Dirac nodal line (black) and the Dirac semimetals (orange), respectively.
@@
Topological Protection in CuMnAs—The altermagnetic surface states discussed above are not topologically protected and might therefore be hard to observe in practice. Imperfect surface termination and disorder can gap them out. In this section, we explain how topologically protected altermagnetic surface states emerge in spin degenerate AFMs. We construct a minimal model that can be tuned to feature altermagnetic drumhead states and Fermi arcs. This minimal model is inspired by the room temperature AFM CuMnAs in its orthorhombic polymorph, reproducing its symmetry properties and band structure features, as we derive rigorously in the SM.
@@
Our starting point is a collinear, inversion odd magnetic order. For simplicity, we consider the same lattice structure and magnetic unit cell as above, see Fig. 1 (d) where the atomic sites represent the 4a Wyckoff position of orthorhombic space group Pnma (SG #62). Hence, the momentum-space Hamiltonian $H_{\text{topo}}=\sum_{\bm{k}}\bm{\Psi}_{\bm{k}}^{\dagger}h(\bm{k})\bm{\Psi}_{\bm{k}}$ and its $8\times 8$ Bloch matrix $h$ are composed of the sublattice space $\sigma$, the layer subspace $\tau$, and spin space $s$. We consider two contributions to the Bloch matrix, $h=h_{0}+h_{M}$. The non-magnetic part of the Hamiltonian,
@@
We consider magnetic order with the same symmetry as that found in CuMnAs, which is inversion odd and hence preserves $\mathcal{T}\parallel\mathcal{TI}]$. In addition, the magnetic order also preserves a glide mirror symmetry combined with time reversal symmetry $\mathcal{T}\parallel\mathcal{TM}_{x}]$. The magnetic order is inversion odd, with the inversion center on each layer. Hence, it is described by spin-dependent hoppings instead of spin-dependent onsite potentials. Note that the spin dependent hoppings are not SOC terms; unlike SOC they are time reversal odd. Considering only up to NNN magnetism-induced hoppings, the magnetic component of $h$, $h_{M}$, has four symmetry-allowed spin-dependent terms
@@
Additionally, the band structure has a topological feature, a Dirac nodal line in the $k_{z}=0$ plane around the $Y$-point, which is visible as a feature in the bulk spectral density, see Fig. 3 (b-c). This four-fold degenerate band crossing is induced by the magnetic hoppings $\Delta_{0},\Delta_{1}$, and protected by a $z$-inverting mirror symmetry. Following the bulk-boundary correspondence, the Dirac nodal line implies the presence of a 2D surface state — a so called drumhead state [4]. This in-gap state is almost flat in energy, but becomes dispersive through the NNN magnetic hoppings $\Delta_{2}$ and $\Delta_{3}$, an is visible in the surface-projected spectral function, see Fig. 3 (d-e).
Deep within the bulk, see Fig. 3 (b-c), the spectral function is spin degenerate. However at the surface, see Fig. 3 (d-e), the drumhead state is partly spin-polarized. All other momentum states have major support in the bulk and are spin degenerate for the same reason as before; a combined global inversion and time reversal symmetry $\mathcal{TI}$ iiiiii Here, the global inversion symmetry only exists exactly if there is an odd number of layers, because the inversion center is on the layer.. The spin polarisaion of the drumhead state on one surface is inverted with respect to the opposing surface. Each drumhead state is protected by weak topology and $d$-wave-like spin-split due to the $\mathcal{T}\parallel\mathcal{TM}_{x}]$ symmetry. The total spectral density at the Fermi level is dominated by the drumhead surface states. We observe that the altermagnetic spin texture leads to a large spin-split conductivity around each surface, see Fig. 3 (j).
The Dirac nodal line is protected by a $z$-inverting mirror symmetry. Breaking this symmetry, e.g. by shear strain, the topological phase changes from a Dirac nodal line semimetal to a Dirac semimetal [36]. In the following, we introduce a term which gaps out the Dirac nodal line into four gapless Dirac points, such that the surface state transforms into four altermagnetic Fermi arcs. This term is an intra-sublattice, intralayer hopping along the (110) direction
@@
These topological Dirac points enforce the formation of altermagnetic Fermi arcs between their projected locations on the surface [22]. The Dirac points consist of a spin-up Weyl point and a spin-down Weyl-point of opposite charge. This implies that each Dirac point must have one spin-up and one spin-down Fermi arc connecting it to an inverted Dirac point. The combined $x$-mirror time reversal symmetry guarantees that the spin up and down Fermi arcs have a different $d$-wave-like momentum dependence. The Fermi arcs can be seen in Fig. 3 (h-i). Considering the slab as a whole, these Fermi arcs are spin degenerate; a Fermi arc on the top surface is identical to its spin-inverted partner on the other side of the system due to global $\mathcal{I}\mathcal{T}$ symmetry. The altermagnetic spin density dominates the density of states, such that the conductivity is strongly spin-split throughout the entire system, see Fig. 3 (k) iiiiiiiiiThe spectral weight in the bulk of the Weyl AM is extremely low since the only gapless states appear around the Dirac points, allowing the edge contributions to dominate throughout the system..
@@
Finally, we also derived the leading order symmetry-allowed SOC terms and studied their effect on the altermagnetic surface states, see SM. Certain SOC terms can lift the topological protection of the altermagnetic surface, however their characteristics remain persistent. Surprisingly, some manifestations of SOC seem to enhance the altermagnetic character of the surface state, leading to states with a much larger spin-split conductivity, see SM. In CuMnAs specifically, SOC is expected to be small, $\approx 7$meV [13].
@@
Conclusion—We have proposed a physical mechanism for topologically protected surface altermagnetism—the formation of robust AM surface states in spin-degenerate AFMs. We first discussed the requirements to realize surface AM: the absence of any combined spin-inversion times in-plane translation symmetries, $C_{2}\parallel\bm{t}_{\text{x or y}}]$, and out-of-plane time reversal two-fold rotation axis, $\mathcal{T}C_{2z}$. These guarantee that the spin-degeneracy is only protected by time-reversal times inversion symmetry, $\mathcal{I}\mathcal{T}$, or spin-inversion times out-of-plane translation symmetry, $C_{2}\parallel\bm{t}_{z}]$. Then, any spin compensated surface is altermagnetic. We illustrated this trivial surface AM in a model where conventional Neél AFM induces a spin split surface.

### v2

@@ Abstract
Here, we take a different path towards realizing AM: on the surface of conventional AFMs. From a symmetry perspective the surface of a crystal is interesting because bulk symmetries are typically broken at the surface by the loss of translational invariance and any point group symmetries incompatible with the surface normal. This symmetry reduction plays a central role in the emergence of surface states with properties forbidden in the bulk, such as in magnetic and topological systems [11, manchon_new_2015]. Furthermore, in many materials, surface states benefit from topological protection, such that they are robust against all symmetry-preserving surface perturbations. This raises the enticing possibility of topologically protecting AM surface states in a conventional AFM—which has yet to be demonstrated and is the focus of this investigation.
@@
As a concrete example for surface magnetism, consider the surface of a conventional AFM. Cutting the crystal along a plane such that only a single spin species exists at the surface, one generates a net surface magnetization—a surface ferromagnet with a spin-split surface state [3, 7, 41]. Similarly, an AFM with non-vanishing surface octupolar moment exhibits an altermagnetization [22, 8]. Although the realization of AM at the surface of an AFM was proposed early on, research has focused so far on identifying a spin-split band structure [39, 56, 40]. However, spin-splitting is not a useful signature for surface AM. The characteristic altermagnetic spin density can emerge as a local feature at the surface. Global bulk properties, e.g., the band structure, cannot detect this effect, because the contributions of opposite surfaces cancel out.
@@
Figure 1: Altermagnetic surface states on spin-degenerate AFMs. (a) The antiferromagnetic bulk symmetry group can break into an altermagnetic surface symmetry group, leading to a spin-split surface Fermi surface (red,blue is the spin character) where the bulk is spin degenerate. (b) In AFM Dirac nodal line semimetals spin-degeneracy is lost at the surface resulting in topologically protected altermagnetic drumhead states. (c) In AFM Dirac semimetals altermagnetic Fermi arcs connect the bulk topological nodes. (d) To unearth the surface behavior described in (a-c) crystals are studied in a slab geometry, i.e., periodic boundary conditions along $x$ and $y$ direction and open boundary conditions in $z$ direction. Here, the crystal structure of Eq. (1) is shown, where spin-up (spin-down) sites are red (blue). The bulk 4-site unit cell is highlighted in blue, labeled by sublattice, and the layer degree of freedom is in the $z$ direction.
@@
In this work, we demonstrate that spin-degenerate AFMs can exhibit altermagnetic spin-split characteristics at the surface. Such a system may be understood as a bulk AFM with local altermagnetic surface states. Note this is substantially different from the surface states of bulk AMs [45, 16, 15, 50, 35, 37, 12, 36, 53] or the influence of an altermagnetic layer on non-magnetic surface states [23]. Of special interest are the cases where these surface states are topologically protected by a bulk symmetry such that are robust.
@@
We report on both trivial and topologically-protected altermagnetic surface states, providing minimal models that realize both of them. First, we explain the general symmetry prerequisites for such surface states to emerge and demonstrate the existence of surface AM in a simple, topologically trivial system with spin-degenerate bulk bands and a spin-split surface, see Fig. 1 (a). Next, we use topology as a resource to protect the altermagnetic surface. Finding such a protected state is challenging because topology in collinear magnets is typically induced by spin-orbit coupling (SOC) [Nagaosa10, Bernevig22], however AM is not well-defined in the presence of SOC [49]. Despite the symmetry conditions for band topology and surface AM being close to incompatible, we find that antiferromagnetic nodal line and Dirac semimetals can exhibit topologically-protected altermagnetic surface states. We find a realization both of AM drumhead states, depicted in Fig. 1 (b), and AM Fermi arcs, depicted in Fig. 1 (c). Finally, we show that the room-temperature antiferromagnetic Dirac semimetals CuMnAs and CuMnP [38], known for their applications in spintronics [51, 47, 46], are existing material realizations of our theory.
@@
Symmetry requirements—Surface AM arises in bulk collinear AFMs when the non-altermagnetic bulk symmetry group breaks down at the surface to an altermagnetic symmetry group. We start by providing the exact symmetry requirements for surface AM effect to occur.
@@
Without loss of generality we assume a (001) surface. Hence, there are two symmetries in a collinear magnet which protect spin-degeneracy in the bulk, but are absent at the surface: (i) an inversion $\mathcal{I}$ times time reversal symmetry $\mathcal{T}\parallel\mathcal{IT}]$ and (ii) a $C_{2}$ rotation in spin space, effectively acting as a spin inversion, times a fractional, (intra-unitcell) translation that involves some translation $\bm{t}_{z}$ in the $z$-direction $C_{2}\parallel\bm{t}_{z}\bm{t}_{\text{any}}]$—here $\bm{t}_{\text{any}}$ indicates that the translation can have any component in the $x$ or $y$ direction. At least one of these two symmetries has to be present to get a bulk AFM, but there must not be any other symmetry that protects spin degeneracy and is robust at the surface, e.g., $C_{2}\parallel\bm{t}_{\text{x or y}}]$. Additionally, there must not be any $\mathcal{T}\parallel C_{2z}\mathcal{T}]$ axis, which would forbid spin splitting at the (001) surface. Finally, one combined spatial and time-reversal symmetry is required to enforce a vanishing magnetic moment at the surface, e.g. $\mathcal{T}\parallel\mathcal{T}C_{4z}]$ and $\mathcal{T}\parallel\mathcal{T}M]$ where $M$ is a mirror. The surface state is then altermagnetic, up to known exceptions [32].
@@
where $m$ is the local magnetisation that leads to the formation of Néel ordering, $s={\uparrow},{\downarrow}$ is the spin, $t=1$ parametrises the nearest neighbour hoppings, $t^{\prime}$ controls the strength of the next-nearest neighbour hoppings within each layer, and $\delta$ determines the degree of anisotropy in the NNN hoppings, see Fig. 1 (d). Here, we consider only this non-interacting effective Hamiltonian, however Eq. (1) may be straightforwardly derived as mean-field description of a Hubbard Hamiltonian with sub-leading NNN couplings, see SM.
@@
Individually, each $xy$-layer of this model would be $d$-wave altermagnetic, because its magnetic sublattices are related by a $\pi/2$-rotation; to this checkerboard square lattice is referred as DLKK model [10, 29, 30] or $tt^{\prime}-$$\delta$-Hubbard model [10, 19]. However, neighboring layers have inverted spin patterns, ensuring that the spin splitting on odd layers perfectly compensates that on even layers. Thus, the overall magnetic state forms a bulk AFM. Formally, the spin up and spin down sublattices are related by a fractional translation $\bm{t}_{z}$, and inversion $\mathcal{I}$, which maps adjacent layers onto one another. In spin-space group notation the symmetries $\mathcal{T}\parallel\mathcal{T}\mathcal{I}]$ and $C_{2}\parallel\bm{t}_{z}]$ protect the spin degeneracy of the band structure, see Fig. 2 (a).
@@
At a (001) surface the $\mathcal{T}\parallel\mathcal{T}\mathcal{I}]$ and $C_{2}\parallel\bm{t}_{z}]$ symmetry are broken. Hence, the surface states in this antiferromagnet are not locally protected against spin splitting. To see this, we consider Eq. 1 in a slab geometry, i.e., a lattice of $L_{z}$ $x$-$y$-layers with open boundary conditions in the $z$-direction and periodic boundary conditions in the $x$- and $y$-directions. Retaining $x$-$y$ translation symmetry allows us to transform to momentum space $\bm{k}=(k_{x},k_{y})$, such that the $L_{z}$ layers appear as $4L_{z}$ different bands in the slab-geometry band structure, see Fig. 2 (b). This model is physically motivated from a Hubbard model—from unrestricted mean-field simulations we found that the magnetization in the 3D slab system is roughly constant throughout the system, see SM.
@@
Surprisingly, even the slab band structure is spin degenerate. This is due to a global inversion times time-reversal symmetry $\mathcal{T}\parallel\mathcal{T}\mathcal{I}]$, which forces momentum states with opposite spin at opposite surfaces of the system to have the same energy iii In fact the global inversion symmetry only exists if there is an even number of layers. For an odd number of layers it is weakly broken but this is an $\frac{1}{L_{z}}$ effect and means that the dispersion of the surface states on opposite sides of the system is slightly different. We expect this splitting in general to be weakly present, but is not the defining spin splitting. This is the splitting which has been identified as surface AM in previous works [39, 56, 40].. However, this has deeper implications: even though spin up and down $\bm{k}$-momentum states are symmetry-constrained to have the same energy they are highly non-local. Due to inversion, their support is at opposite sides of the system. Hence, we can observe a spin-split surface state on one side of the system and its spin-inverted counterpart on the other side of the system.
Figure 2: Topologically trivial altermagnetic surface state in a spin degenerate AFM, defined in Eq. 1. (a) The 3D bulk band structure is spin-degenerate. (b-e) Colorplots of the spin-resolved spectral function with the following color scheme: $\mathcal{A}_{{\uparrow}}-\mathcal{A}_{{\downarrow}}$ sets the coloring from spin-down polarized (blue) over spin-degenerate (black) to spin-up polarized (red); $\mathcal{A}_{{\uparrow}}+\mathcal{A}_{{\downarrow}}$ sets the opacity. (b,c) The bulk spin spectral density $\mathcal{A}_{s}(z=L_{z}/2)$ in the slab geometry ($L_{z}=30$) along a high symmetry path (c) and at the Fermi energy ($E=0$, dashed line) (d). (d,e) The spin spectral density $\mathcal{A}_{s}(z=0)$ in the slab geometry at the surface along a high symmetry path (c) and at the Fermi energy (d). (f) The layer-resolved spin conductivity, expressed as the spin splitter angle $\alpha=2\arctan((\sigma^{\uparrow}_{xx}-\sigma^{\downarrow}_{xx})/(\sigma^{\uparrow}_{xx}+\sigma^{\downarrow}_{xx}))$ shows that the the spin splitting is only sizable at the surface.
@@
We visualize the spin-, momentum-, and space-resolved spectral density by a spin-resolved spectral function, which may be written in terms of the Greens function, $G(\omega,\bm{k})$ as
@@
Finally, we also computed the layer-resolved (again we consider pairs of layers) spin conductance, see SM for details. Here, we use the spin-splitter angle $\alpha$ to quantify the spin-splitting, i.e., the angle between the spin-up and down current which is determined via the ratio of spin over electric conductivity [17]. As expected from the spin-split surface state, the spin conductance is only non-zero at the surface and odd under global inversion, see Fig. 2 (f).
@@
Figure 3: Topologically protected altermagnetic surface states in an AFM, described by $H_{\text{topo}}$ following Eqs. 3 and 4. (a) The 3D bulk band structure of the Dirac nodal line semimetal (black, $t^{\prime}=0$) and the Dirac semimetal (yellow, $t^{\prime}=0.2$) is spin degenerate. Dashed lines indicate overlap of both band structures. The touching points at E=0.1E=-0.1 (shifted with a small chemical potential $\mu$) are the nodal line and the Dirac points. The spin-resolved spectral function in a slab geometry ($L_{z}=31$) of the Dirac nodal line semimetal is shown in (b-e) and the Dirac semimetal in (f-i). The color scheme is identical to Fig. 2. Panel (b,f) and (d,h) show the bulk and surface spin spectral density of each model along a high symmetry path respectively. The bulk spin spectral density $\mathcal{A}_{s}(z=15)$ at the Fermi energy ($E=0$, dashed line) shows the Dirac nodal line (c) and the Dirac cones (g). The surface spin spectral density $\mathcal{A}_{s}(z=0)$ at the Fermi energy shows the corresponding surface states, where closed contours are crossings of the drumhead states and the Fermi energy (e) and open arcs are crossings of the Fermi arcs and the Fermi energy (i). (j,k) The layer-resolved spin-splitter angles $\alpha_{x}$ ($\alpha_{y}$), as solid (dashed) line, which can be observed when applying an electric field in $x$-direction ($y$-direction). Panels (j) and (k) correspond to the Dirac nodal line (black) and the Dirac semimetals (orange), respectively.
@@
Topological protection in CuMnAs—The altermagnetic surface states discussed above are not topologically protected and therefore be hard to observe in practice. Imperfect surface termination and disorder can gap them out. In this section, we explain how topologically protected altermagnetic surface states emerge in spin degenerate AFMs. We construct a minimal model that can be tuned to feature altermagnetic drumhead states and Fermi arcs. This minimal model is inspired by the room temperature AFM CuMnAs in its orthorhombic polymorph, reproducing their symmetry properties and band structure features, as we derive rigorously in the SM.
@@
Our starting point is a collinear, inversion odd magnetic order. For simplicity, we consider the same lattice structure and magnetic unit cell as above, see Fig. 1 (d), where the atomic sites represent the 4a Wyckoff position of orthorhombic space group Pnma (SG #62). Hence, the momentum-space Hamiltonian $H_{\text{topo}}=\sum_{\bm{k}}\bm{\Psi}_{\bm{k}}^{\dagger}h(\bm{k})\bm{\Psi}_{\bm{k}}$ and its $8\times 8$ Bloch matrix $h$ are composed of the sublattice space $\sigma$, the layer subspace $\tau$, and spin space $s$. We consider two contributions to the Bloch matrix, $h=h_{0}+h_{M}$. The non-magnetic part of the Hamiltonian,
@@
We consider a magnetic order with the same symmetry as found in CuMnAs, which is inversion odd and hence preserves $\mathcal{T}\parallel\mathcal{TI}]$. In addition, the magnetic order also preserves a glide mirror symmetry combined with time reversal symmetry $\mathcal{T}\parallel\mathcal{TM}_{x}]$. Since the sites are the inversion centers, the magnetic order is described by spin-dependent hoppings instead of spin-dependent onsite potentials. Note that these hoppings are not SOC terms; unlike SOC they are time reversal odd. Considering only up to NNN magnetism-induced hoppings, the magnetic component of $h$, $h_{M}$, has four symmetry-allowed spin-dependent terms
@@
Additionally, the band structure has a topological feature, a Dirac nodal line in the $k_{z}=0$ plane around the $Y$-point, which is visible as a feature in the bulk spectral density, see Fig. 3 (b-c). This four-fold degenerate band crossing is induced by the magnetic hoppings $\Delta_{0},\Delta_{1}$ and protected by a $z$-inverting mirror symmetry. Following the bulk-boundary correspondence, the Dirac nodal line implies the presence of a 2D surface state — a so called drumhead state [4]. This in-gap state is almost flat in energy, but becomes dispersive through the NNN magnetic hoppings $\Delta_{3}$ (or $\Delta_{2}$, we set $\Delta_{2}=0$), and is visible in the surface-projected spectral function, see Fig. 3 (d-e).
Deep within the bulk, see Fig. 3 (b-c), the spectral function is spin degenerate iiiiii Here, the global inversion symmetry only exists exactly if there is an odd number of layers, because the inversion center is on the layer.. However at the surface, see Fig. 3 (d-e), the drumhead state is partly spin-polarized. The spin polarization of the drumhead state on one surface is inverted with respect to the opposing surface. Each drumhead state is protected by weak topology and $d$-wave-like spin-split due to the $\mathcal{T}\parallel\mathcal{TM}_{x}]$ symmetry. The total spectral density at the Fermi level is dominated by the drumhead surface states. We observe that the altermagnetic spin texture leads to a large spin-split conductivity around each surface, see Fig. 3 (j).
The Dirac nodal line is protected by a $z$-inverting mirror symmetry. By breaking this symmetry, e.g. by shear strain, the topological phase changes from a Dirac nodal line semimetal to a Dirac semimetal [38, 6]. In the following, we introduce a term which gaps out the nodal line into four gapless Dirac points, such that the surface state transforms into four altermagnetic Fermi arcs. This term is an intrasublattice, intralayer hopping along the (110) direction
@@
The topological Dirac points enforce the formation of altermagnetic Fermi arcs between their projected locations on the surface [24]. The Dirac points consist of a spin-up Weyl point and a spin-down Weyl-point of opposite charge. This implies that each Dirac point must have one spin-up and one spin-down Fermi arc connecting it to an inverted Dirac point. The combined $x$-mirror time reversal symmetry guarantees that the spin up and down Fermi arcs have a different $d$-wave-like momentum dependence, see Fig. 3 (h-i), with the opposite surface having an inverted spin configuration. The altermagnetic spin density dominates the density of states, such that the conductivity is strongly spin-split throughout the entire system, see Fig. 3 (k) iiiiiiiiiThe spectral weight in the bulk of the Weyl AM is extremely low since the only gapless states appear around the Dirac points, allowing the edge contributions to dominate throughout the system..
@@
Finally, we also derived the leading order symmetry-allowed SOC terms and studied their effect on the altermagnetic surface states, see SM. Certain SOC terms can lift the topological protection of the altermagnetic surface, however its characteristics remain persistent. Surprisingly, some manifestations of SOC seem to enhance the altermagnetic character of the surface state, leading to a larger spin-split conductivity, see SM. In CuMnAs specifically, SOC is expected to be small, $\approx 7$meV [14].
@@
Conclusion—We have proposed a physical mechanism for topologically protected surface altermagnetism—the formation of robust AM surface states in spin-degenerate AFMs. We first discussed the requirements to realize surface AM: the absence of any combined spin-inversion times in-plane translation symmetries, $C_{2}\parallel\bm{t}_{\text{x or y}}]$, and out-of-plane time reversal two-fold rotation axis, $\mathcal{T}C_{2z}$. These guarantee that the spin-degeneracy is only protected by time-reversal times inversion symmetry, $\mathcal{I}\mathcal{T}$, or spin-inversion times out-of-plane translation symmetry, $C_{2}\parallel\bm{t}_{z}]$. Then, any spin compensated surface is altermagnetic. We illustrated a resulting trivial surface AM in a model where conventional Neél AFM induces a spin split surface.